OpenStudy (anonymous):

The surface area of the top surface of the water in a circular swimming pool is about 206 square feet. Estimate the radius of the pool, to the nearest foot.

OpenStudy (unklerhaukus):

\[A=\pi R^2\] \[\sqrt{\frac{A}{\pi}}=R\]

OpenStudy (anonymous):

so do you take 260/ 3.14?

OpenStudy (unklerhaukus):

260/ 3.14~86 now the square root of that, is a bit more than the square root of 81 which is nine

OpenStudy (anonymous):

there is no answer of 9 gosh i dont get this at all

OpenStudy (unklerhaukus):

ops the area is 206, not 260

OpenStudy (anonymous):

ooops

OpenStudy (unklerhaukus):

(206/3)~68.7 the square root of that is a bit more than the square root of 64 ie 8

OpenStudy (unklerhaukus):

so the pool has a radius of a little bit more than 8ft

OpenStudy (anonymous):

ty:)

OpenStudy (anonymous):

Estimate the square root to the nearest integer. of 25

OpenStudy (anonymous):

thats another one im stuck on